OpenStudy (anonymous):

What are the x- and y-intercepts of the quadratic equation y = (x - 1)2 - 16? Identify the x-intercepts by graphing the quadratic equation.

OpenStudy (campbell_st):

well start by expanding and collecting like terms \[y = x^2 - 2x + 1 - 16\] so \[y = x^2 -2x - 15\] the y intercept occurs when x = 0... so substitute x = 0 and evaluate the x intercept is who y = 0 so you need to solve \[0=x^2 -2x - 15\] which can be factored. hope this helps